OpenStudy (anonymous):

how to factor 9x^2-16

OpenStudy (cwrw238):

this is difference of 2 squares use the identity a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

its confusing me because the x^2 is part of the nine

OpenStudy (cwrw238):

a^2 is equavalent to 9x^2 so a -------------- 3x right?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

and b^2 equivalent to 16 b …………………………..?

OpenStudy (anonymous):

4

OpenStudy (skullpatrol):

To factor a difference of two squares you need to recognize that the exponents of all the powers in it are even and the numerical coefficient is the square of an integer.

OpenStudy (cwrw238):

right so plug those values into RHS of the identity

OpenStudy (anonymous):

so its 3^2-4^2?

OpenStudy (skullpatrol):

$$\huge 9x^2-16=(3x)^2-4^2$$

OpenStudy (cwrw238):

no just plug in a = 3x and b = 4 into (a + b)(a - b)

OpenStudy (anonymous):

Oh i get it so its (3x+4)(3x-4)

OpenStudy (cwrw238):

yes

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

but what if its x^3+27? how would I go about solving that?

OpenStudy (skullpatrol):

That is a sum of cubes.

OpenStudy (anonymous):

do i do (x2+9)(x+3)

OpenStudy (cwrw238):

just look out for difference of 2 perfect squares eg x^ 2 - 9 16y^2 - 25

OpenStudy (skullpatrol):

$$\huge x^3+27=x^3+3^3$$

OpenStudy (anonymous):

for the second equation the answer is (4y-5)(4y+5) :)

OpenStudy (anonymous):

skullpatrol how do i put that in the equation cwrw was showing me?

OpenStudy (skullpatrol):

It requires a different identity.

OpenStudy (anonymous):

what is that identity?

OpenStudy (skullpatrol):

$$\huge x^3+y^3=(x+y)(x^2-xy+y^2)$$

OpenStudy (anonymous):

how do i plug it in then?

OpenStudy (skullpatrol):

$$\huge x^3+3^3$$

OpenStudy (anonymous):

do i have to put it in ( ) ( ) that form or can i just leave my answer as x^3+3^3

OpenStudy (skullpatrol):

You have to put it in ( ) ( ) form.

OpenStudy (anonymous):

how do i do that? im sorry im very slow especially when it comes to math :/

OpenStudy (skullpatrol):

Use the identity with x = x and y = 3.

OpenStudy (anonymous):

so it'll look like (x+3)(x^2-x3+3^2) ?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

but that makes no sense

OpenStudy (skullpatrol):

why

OpenStudy (anonymous):

because if yo solve it it doesnt come out to be x^3+3^3

OpenStudy (anonymous):

its okay I'll just ask my teacher tomorrow Im probably frustrating you now thanks for the help though!

OpenStudy (skullpatrol):